OpenStudy (captain_deadpool2653):

i need help on reducing a fraction k+7 --- 4k+28

OpenStudy (photon336):

okay so at the bottom you've got 4k+28 if you notice, we can factor out a 4

OpenStudy (captain_deadpool2653):

ok

OpenStudy (photon336):

So the bottom becomes this \[4(k+7)\] can you see why?

OpenStudy (captain_deadpool2653):

yes

OpenStudy (photon336):

if you multiply 4*k+4*7 you get 4k+28

OpenStudy (photon336):

now here's what we have \[\frac{ (k+7) }{ 4(k+7) }\]

OpenStudy (photon336):

so @Captain_Deadpool2653 because (k+7) is in both the numerator and denominator, you end up getting \[\frac{ (\cancel\k+\cancel\7) }{ 4 \cancel\k+\cancel\7 } = \frac{ 1 }{ 4 }\]

OpenStudy (captain_deadpool2653):

thanks that really helped

OpenStudy (photon336):

no problem